LILO

Generic boot loader for Linux Version 20

User’s guide

Werner Almesberger

                   werner.almesberger@lrc.di.epfl.ch

June 20, 1997

Contents

1 Introduction 4 1.1 Disk organization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Booting basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3 Choosing the “right” boot concept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2 The boot prompt 14 2.1 Boot command-line options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.2 Boot image selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 Map installer 18 3.1 Command-line options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.2 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.3 Boot image types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.4 Disk geometry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.5 Partition table manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.6 Keyboard translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

4 Installation and updates 33 4.1 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2 Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.3 LILO de-installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.4 Installation of other operating systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

5 Troubleshooting 40 5.1 Map installer warnings and errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.2 Boot loader messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.3 Other problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

LILO is a versatile boot loader for Linux. It does not depend on a specific file system, can boot Linux kernel images from floppy disks and from hard disks and can even act as a “boot manager” for other operating systems.^{1}

One of up to sixteen different images can be selected at boot time. Various parameters, such as the root device, can be set independently for each kernel. LILO can even be used as the master boot record.

1PC/MS-DOS, DR DOS, OS/2, Windows 95, Windows NT, 386BSD, SCO UNIX, Unixware, …

This document introduces the basics of disk organization and booting, continues with an overview of common boot techniques and finally describes installation and use of LILO in greater detail. The troubleshooting section at the end describes diagnostic messages and contains suggestions for most problems that have been observed in the past.

Please read at least the sections about installation and configuration if you’re already using an older version of LILO. This distribution is accompanied by a file named INCOMPAT that describes further incompatibilities to older versions.

For the impatient: there is a quick-installation script to create a simple but quite usable installation. See section 4.1.2 for details.

But wait … here are a few easy rules that will help you to avoid most problems people experience with LILO:

• Don’t panic. If something doesn’t work, try to find out what is wrong, try to verify your assumption and only then attempt to fix it.

• Read the documentation. Especially if what the system does doesn’t correspond to what you think it should do.

• Make sure you have an emergency boot disk, that you know how to use it, and that it is always kept up to date.

• Run /sbin/lilo whenever the kernel or any part of LILO, including its configuration file, has changed. When in doubt, run it. You can’t run /sbin/lilo too many times.

• If using LILO as the MBR, de-install it before performing a destructive upgrade and/or erasing your Linux partitions.

• Don’t trust setup scripts. Always verify the /etc/lilo.conf they create before booting.

• If using a big disk, be prepared for inconveniences: you may have to use the linear option.

System overview

LILO is a collection of several programs and other files:

The map installer is the program you run under Linux to put all files belonging to LILO at the appropriate places and to record information about the location of data needed at boot time. This program normally resides in /sbin/lilo. It has to be run to refresh that information whenever any part of the system that LILO knows about changes, e.g. after installing a new kernel.

Various files contain data LILO needs at boot time, e.g. the boot loader. Those files normally reside in /boot. The most important files are the boot loader (see below) and the map file (/boot/map), where the map installer records the location of the kernel(s).^{2} Another important file is the configuration file, which is normally called /etc/lilo.conf

The boot loader is the part of LILO that is loaded by the BIOS and that loads kernels or the boot sectors of other operating systems. It also provides a simple command-line interface to interactively select the item to boot and to add boot options.

LILO primarily accesses the following parts of the system:

The root file system partition is important for two reasons: first, LILO sometimes has to tell the kernel where to look for it. Second, it is frequently a convenient place for many other items LILO uses, such as the boot sector, the /boot directory, and the kernels.

The boot sector contains the first part of LILO’s boot loader. It loads the much larger second-stage loader. Both loaders are typically stored in the file /boot/boot.b

The kernel is loaded and started by the boot loader. Kernels typically reside in the root directory or in

/boot.

Note that many of the files LILO needs at boot time have to be accessible with the BIOS. This creates certain restrictions, see section 1.3.1.

2LILO does not know how to read a file system. Instead, the map installer asks the kernel for the physical location of files (e.g. the kernel image(s)) and records that information. This allows LILO to work with most file systems that are supported by Linux.

1 Introduction

The following sections describe how PCs boot in general and what has to be known when booting Linux and using LILO in particular.

1.1 Disk organization

When designing a boot concept, it is important to understand some of the subtleties of how PCs typically organize disks. The most simple case are floppy disks. They consist of a boot sector, some administrative data (FAT or super block, etc.) and the data area. Because that administrative data is irrelevant as far as booting is concerned, it is regarded as part of the data area for simplicity.

Boot sector Data area

The entire disk appears as one device (e.g. /dev/fd0) on Linux.

The MS-DOS boot sector has the following structure:

0x000 Jump to the program code

                                 0x003

Disk parameters

                           0x02C/0x03E

Program code

0x1FE Magic number (0xAA55)

LILO uses a similar boot sector, but it does not contain the disk parameters part. This is no problem for Minix, Ext2 or similar file systems, because they don’t look at the boot sector, but putting a LILO boot sector on an MS-DOS file system would make it inaccessible for MS-DOS.

Hard disks are organized in a more complex way than floppy disks. They contain several data areas called partitions. Up to four so-called primary partitions can exist on an MS-DOS hard disk. If more partitions are needed, one primary partition is used as an extended partition that contains several logical partitions.

The first sector of each hard disk contains a partition table, and an extended partition and each logical partition contains a partition table too.

Partition table /dev/hda Partition 1 /dev/hda1

Partition 2 /dev/hda2

The entire disk can be accessed as /dev/hda, /dev/hdb, /dev/sda, etc. The primary partitions are

/dev/hda1 . . . /dev/hda4.

Partition table /dev/hda Partition 1 /dev/hda1

Partition 2 /dev/hda2

Extended partition /dev/hda3 Extended partition table Partition 3 /dev/hda5

Extended partition table

Partition 4 /dev/hda6

This hard disk has two primary partitions and an extended partition that contains two logical partitions. They are accessed as /dev/hda5 …

Note that the partition tables of logical partitions are not accessible as the first blocks of some devices, while the main partition table, all boot sectors and the partition tables of extended partitions are.

Partition tables are stored in partition boot sectors. Normally, only the partition boot sector of the entire disk is used as a boot sector. It is also frequently called the master boot record (MBR). Its structure is as follows:

                               0x000

Program code

0x1BE  Partition table
0x1FE Magic number (0xAA55)

The LILO boot sector is designed to be usable as a partition boot sector. (I.e. there is room for the partition table.) Therefore, the LILO boot sector can be stored at the following locations:

• boot sector of a Linux floppy disk. (/dev/fd0, …)

• MBR of the first hard disk. (/dev/hda, /dev/sda, …)

• boot sector of a primary Linux file system partition on the first hard disk. (/dev/hda1, …)

• partition boot sector of an extended partition on the first hard disk. (/dev/hda1, …)^{3}

It can’t be stored at any of the following locations:

• boot sector of a non-Linux floppy disk or primary partition.

• a Linux swap partition.

• boot sector of a logical partition in an extended partition.^{4}

• on the second hard disk. (Unless for backup installations, if the current first disk will be removed or disabled, or if some other boot loader is used, that is capable of loading boot sectors from other drives.)

3Most FDISK-type programs don’t believe in booting from an extended partition and refuse to activate it. LILO is accompanied by a simple program (activate) that doesn’t have this restriction. Linux fdisk also supports activating extended partitions.

4LILO can be forced to put the boot sector on such a partition by using the -b option or the boot variable. However, only few programs that operate as master boot records support booting from a logical partition.

Although LILO tries to detect attempts to put its boot sector at an invalid location, you should not rely on that.

1.2 Booting basics

When booting from a floppy disk, the first sector of the disk, the so-called boot sector, is loaded. That boot sector contains a small program that loads the respective operating system. MS-DOS boot sectors also contain a data area, where disk and file system parameters (cluster size, number of sectors, number of heads, etc.) are stored.

When booting from a hard disk, the very first sector of that disk, the so-called master boot record (MBR) is loaded. This sector contains a loader program and the partition table of the disk. The loader program usually loads the boot sector, as if the system was booting from a floppy.

Note that there is no functional difference between the MBR and the boot sector other than that the MBR contains the partition information but doesn’t contain any file system-specific information (e.g. MS-DOS disk parameters).

The first 446 (0x1BE) bytes of the MBR are used by the loader program. They are followed by the partition table, with a length of 64 (0x40) bytes. The last two bytes contain a magic number that is sometimes used to verify that a given sector really is a boot sector.

There is a large number of possible boot configurations. The most common ones are described in the following sections.

1.2.1 MS-DOS alone

Master Boot Record Boot sector

Operating system

DOS-MBR ⟶ MS-DOS ⟶ COMMAND.COM

This is what usually happens when MS-DOS boots from a hard disk: the DOS-MBR determines the active partition and loads the MS-DOS boot sector. This boot sector loads MS-DOS and finally passes control to COMMAND.COM. (This is greatly simplified.)

1.2.2 LOADLIN

Master Boot Record Boot sector

Operating systems

DOS-MBR ⟶ MS-DOS ⟶ COMMAND.COM
⟶ LOADLIN ⟶ Linux

A typical LOADLIN setup: everything happens like when booting MS-DOS, but in CONFIG.SYS or AUTOEXEC.BAT, LOADLIN is invoked. Typically, a program like BOOT.SYS is used to choose among configuration sections in CONFIG.SYS and AUTOEXEC.BAT. This approach has the pleasant property that no boot sectors have to be altered.

Please refer to the documentation accompanying the package for installation instructions and further details.

1.2.3 LILO started by DOS-MBR

Master Boot Record Boot sector

Operating system

DOS-MBR ⟶ LILO ⟶ Linux
⟶ other OS

This is a “safe” LILO setup: LILO is booted by the DOS-MBR. No other boot sectors have to be touched. If the other OS (or one of them, if there are several other operating systems being used) should be booted without using LILO, the other partition has to be marked “active” with fdisk or activate.

Installation:

• install LILO with its boot sector on the Linux partition.

• use fdisk or activate to make that partition active.

• reboot. Deinstallation:

• make a different partition active.

• install whatever should replace LILO and/or Linux.

1.2.4 Several alternate branches Master Boot Record Boot sector

Operating systems

DOS-MBR ⟶ MS-DOS ⟶ COMMAND.COM
⟶ LOADLIN ⟶ Linux
⟶ LILO ⟶ Linux
⟶ MS-DOS — ⋯

An extended form of the above setup: the MBR is not changed and both branches can either boot Linux or MS-DOS. (LILO could also boot any other operating system.)

1.2.5 LILO started by BOOTACTV^{5} Master Boot Record Boot sector

Operating system

BOOTACTV ⟶ LILO ⟶ Linux
⟶ other OS

Here, the MBR is replaced by BOOTACTV (or any other interactive boot partition selector) and the choice between Linux and the other operating system(s) can be made at boot time. This approach should be used if LILO fails to boot the other operating system(s).^{6} Installation:

• boot Linux.

• make a backup copy of your MBR on a floppy disk, e.g.

     dd if=/dev/hda of=/fd/MBR bs=512 count=1

• install LILO with the boot sector on the Linux partition.

• install BOOTACTV as the MBR, e.g.

     dd if=bootactv.bin of=/dev/hda bs=446 count=1

• reboot. Deinstallation:

• boot Linux.

5Other, possibly better known boot switchers, e.g. OS/2 BootManager operate in a similar way. The installation procedures typically vary.

6And the author would like to be notified if booting the other operating system(s) doesn’t work with LILO, but if it works with an other boot partition selector.

• restore the old MBR, e.g.

     dd if=/MBR of=/dev/hda bs=446 count=1

or FDISK /MBR under MS-DOS.

If replacing the MBR appears undesirable and if a second Linux partition exists (e.g. /usr, not a swap partition), BOOTACTV can be merged with the partition table and stored as the “boot sector” of that partition. Then, the partition can be marked active to be booted by the DOS-MBR.

Example:

# dd if=/dev/hda of=/dev/hda3 bs=512 count=1
# dd if=bootactv.bin of=/dev/hda3 bs=446 count=1

WARNING: Whenever the disk is re-partitioned, the merged boot sector on that “spare” Linux partition has to be updated too.

1.2.6 LILO alone

Master Boot Record Operating system

LILO ⟶ Linux
⟶ other OS

LILO can also take over the entire boot procedure. If installed as the MBR, LILO is responsible for either booting Linux or any other OS. This approach has the disadvantage, that the old MBR is overwritten and has to be restored (either from a backup copy, with FDISK /MBR on recent versions of MS-DOS or by overwriting it with something like BOOTACTV) if Linux should ever be removed from the system.

You should verify that LILO is able to boot your other operating system(s) before relying on this method.

Installation:

• boot Linux.

• make a backup copy of your MBR on a floppy disk, e.g.

     dd if=/dev/hda of=/fd/MBR bs=512 count=1

• install LILO with its boot sector as the MBR.

• reboot.

Deinstallation:

• boot Linux.

• restore the old MBR, e.g.

     dd if=/fd/MBR of=/dev/hda bs=446 count=1

If you’ve installed LILO as the master boot record, you have to explicitly specify the boot sector (configuration variable boot=…) when updating the map. Otherwise, it will try to use the boot sector of your current root partition, which will usually work, but it will probably leave your system unbootable.

1.2.7 Names

The following names have been used to describe boot sectors or parts of operating systems:

“DOS-MBR” is the original MS-DOS MBR. It scans the partition table for a partition that is marked “active” and loads the boot sector of that partition. Programs like MS-DOS’ FDISK, Linux fdisk or activate (accompanies LILO) can change the active marker in the partition table.

“MS-DOS” denotes the MS-DOS boot sector that loads the other parts of the system (IO.SYS, etc.).

“COMMAND.COM” is the standard command interpreter of MS-DOS.

“LOADLIN” is a program that loads a Linux kernel image from an MS-DOS partition into memory and executes it. It is usually invoked from CONFIG.SYS and is used in combination with a CONFIG.SYS configuration switcher, like BOOT.SYS.^{7}

“LILO” can either load a Linux kernel or the boot sector of any other operating system. It has a first stage boot sector that loads the remaining parts of LILO from various locations.^{8}

“BOOTACTV” permits interactive selection of the partition from which the boot sector should be read. If no key is pressed within a given interval, the partition marked active is booted. BOOTACTV is included in the pfdisk package. There are also several similar programs, like PBOOT and OS-BS.^{9}

1.3 Choosing the “right” boot concept

Although LILO can be installed in many different ways, the choice is usually limited by the present setup and therefore, typically only a small number of configurations which fit naturally into an existing system remains. The following sections describe various possible cases. See also section 1.3.1.

The configuration file /etc/lilo.conf for the examples could look like this:

boot = /dev/hda2
compact
image = /vmlinuz
image = /vmlinuz.old
other = /dev/hda1
  table = /dev/hda
  label = msdos

It installs a Linux kernel image (/vmlinuz), an alternate Linux kernel image (/vmlinuz.old) and a chain loader to boot MS-DOS from /dev/hda1. The option compact on the second line instructs the map installer to optimize loading.

In all examples, the names of the IDE-type hard disk devices (/dev/hda…) are used. Everything applies to other disk types (e.g. SCSI disks; /dev/sda…) too.

1.3.1 BIOS restrictions

Nowadays, an increasing number of systems is equipped with comparably large disks or even with multiple disks. At the time the disk interface of the standard PC BIOS has been designed (about 16 years ago), such configurations were apparently considered to be too unlikely to be worth supporting.

7LOADLIN is available for anonymous FTP from ftp://tsx-11.mit.edu/pub/linux/dos utils/lodlinn.tar.gz ftp://sunsite.unc.edu/pub/Linux/system/boot/lodlinn.tgz BOOT.SYS is available for anonymous FTP from ftp://sunsite.unc.edu/pub/Linux/system/Linux-boot/boot142.zip or ftp://ftp.funet.fi/pub/Linux/tools/boot142.zip

8LILO can be found in ftp://tsx-11.mit.edu/pub/linux/packages/lilo/lilo-n.tar.gz ftp://sunsite.unc.edu/pub/Linux/system/boot/lilo/lilo-n.tar.gz ftp://lrcftp.epfl.ch/pub/linux/local/lilo/lilo-n.tar.gz

9pfdisk is available for anonymous FTP from ftp://sunsite.unc.edu/pub/Linux/utils/disk-management/pfdisk.tar.Z or ftp://ftp.funet.fi/pub/Linux/tools/pfdisk.tar.Z PBOOT can be found in ftp://ftp.funet.fi/pub/Linux/tools/pboot.zip

The most common BIOS restrictions that affect LILO are the limitation to two hard disks and the inability to access more than 1024 cylinders per disk. LILO can detect both conditions, but in order to work around the underlying problems, manual intervention is necessary.

The drive limit does not exist in every BIOS. Some modern motherboards and disk controllers are equipped with a BIOS that supports more (typically four) disk drives. When attempting to access the third, fourth, etc. drive, LILO prints a warning message but continues. Unless the BIOS really supports more than two drives, the system will not be able to boot in that case.^{10}

The cylinder limit is a very common problem with IDE disks. There, the number of cylinders may already exceed 1024 if the drive has a capacity of more than 504 MB. Many SCSI driver BIOSes present the disk geometry in a way that makes the limit occur near 1 GB. Modern disk controllers may even push the limit up to about 8 GB. All cylinders beyond the 1024th are inaccessible for the BIOS. LILO detects this problem and aborts the installation (unless the linear option is used, see section 3.2.2).

Note that large partitions that only partially extend into the “forbidden zone” are still in jeopardy even if they appear to work at first, because the file system does not know about the restrictions and may allocate disk space from the area beyond the 1024th cylinder when installing new kernels. LILO therefore prints a warning message but continues as long as no imminent danger exists.

There are four approaches of how such problems can be solved:

• use of a different partition which is on an accessible disk and which does not exceed the 1024 cylinder limit. If there is only a DOS partition which fulfills all the criteria, that partition can be used to store the relevant files. (See section 1.3.8.)

• rearranging partitions and disks. This is typically a destructive operation, so extra care should be taken to make good backups.

• if the system is running DOS or Windows 95, LOADLIN can be used instead of LILO.

• if all else fails, installation of a more capable BIOS, a different controller or a different disk configuration.

LILO depends on the BIOS to load the following items:

   • /boot/boot.b

• /boot/map (created when running /sbin/lilo)

• all kernels

• the boot sectors of all other operating systems it boots

• the startup message, if one has been defined

Normally, this implies that the Linux root file system should be in the “safe” area. However, it is already sufficient to put all kernels into /boot and to either mount a “good” partition on /boot or to let /boot be a symbolic link pointing to or into such a partition.

See also /usr/src/linux/Documentation/ide.txt (or /usr/src/linux/drivers/block/README.ide in

older kernels) for a detailed description of problems with large disks.

1.3.2 One disk, Linux on a primary partition

If at least one primary partition of the first hard disk is used as a Linux file system (/, /usr, etc. but not as a swap partition), the LILO boot sector should be stored on that partition and it should be booted by the original master boot record or by a program like BOOTACTV.

10However, if only “unimportant” parts of the system are located on the “high” drives, some functionality may be available.

                                    MBR  /dev/hda
                                      MS-DOS /dev/hda1

                               →  Linux /     /dev/hda2

In this example, the boot variable could be omitted, because the boot sector is on the root partition.

1.3.3 One disk, Linux on a logical partition

If no primary partition is available for Linux, but at least one logical partition of an extended partition on the first hard disk contains a Linux file system, the LILO boot sector should be stored in the partition sector of the extended partition and it should be booted by the original master boot record or by a program like BOOTACTV.

                                    MBR  /dev/hda
                                      MS-DOS /dev/hda1

                               →  Extended /dev/hda2

                                        Linux  /dev/hda5
                                        . . .  /dev/hda6

Because many disk partitioning programs refuse to make an extended partition (in our example /dev/hda2) active, you might have to use activate, which comes with the LILO distribution.

OS/2 BootManager should be able to boot LILO boot sectors from logical partitions. The installation on the extended partition itself is not necessary in this case.

1.3.4 Two disks, Linux (at least partially) on the first disk

This case is equivalent to the configurations where only one disk is in the system. The Linux boot sector resides on the first hard disk and the second disk is used later in the boot process.

Only the location of the boot sector matters – everything else (/boot/boot.b, /boot/map, the root file system, a swap partition, other Linux file systems, etc.) can be located anywhere on the second disk, provided that the constraints described in section 1.3.1 are met.

1.3.5 Two disks, Linux on second disk, first disk has an extended partition

If there is no Linux partition on the first disk, but there is an extended partition, the LILO boot sector can be stored in the partition sector of the extended partition and it should be booted by the original master boot record or by a program like BOOTACTV.

First disk  Second disk

                      MBR  /dev/hda  MBR  /dev/hdb
                        MS-DOS /dev/hda1  Linux  /dev/hdb1
                  →  Extended /dev/hda2  . . .  /dev/hdb2
                           . . .  /dev/hda5
                           . . .  /dev/hda6

The program activate, that accompanies LILO, may have to be used to set the active marker on an extended partition, because MS-DOS’ FDISK and some older version of Linux fdisk refuse to do that. (Which is generally a good idea.)

1.3.6 Two disks, Linux on second disk, first disk has no extended partition

If there is neither a Linux partition nor an extended partition on the first disk, then there’s only one place left, where a LILO boot sector could be stored: the master boot record.

In this configuration, LILO is responsible for booting all other operating systems too.

First disk  Second disk

                  → MBR  /dev/hda  MBR  /dev/hdb
                        MS-DOS /dev/hda1  Linux  /dev/hdb1
                        . . .  /dev/hda2  . . .  /dev/hdb2

You should back up your old MBR before installing LILO and verify that LILO is able to boot your other operating system(s) before relying on this approach.

The line boot = /dev/hda2 in /etc/lilo.conf would have to be changed to boot = /dev/hda in this

example.

1.3.7 More than two disks

On systems with more than two disks, typically only the first two can be accessed. The configuration choices are therefore the same as with two disks.

When attempting to access one of the extra disks, LILO displays a warning message (Warning: BIOS drive 0xnumber may not be accessible) but does not abort. This is done in order to allow the lucky few whose BIOS (or controller-BIOS) does support more than two drives to make use of this feature. By all others, this warning should be considered a fatal error.

Note that the two disks restriction is only imposed by the BIOS. Linux normally has no problems using all disks once it is booted.

1.3.8 /boot on a DOS partition

Recent kernels support all the functions LILO needs to map files also on MS-DOS (or UMSDOS) file systems. Since DOS partitions tend to occupy exactly the places where BIOS restrictions (see section 1.3.1) are invisible, they’re an ideal location for /boot if the native Linux file systems can’t be used because of BIOS problems.

In order to accomplish this, the DOS partition is mounted read-write, a directory (e.g. /dos/linux) is created, all files from /boot are moved to that directory, /boot is replaced by a symbolic link to it, the kernels are also moved to the new directory, their new location is recorded in /etc/lilo.conf, and finally

/sbin/lilo is run.

From then on, new kernels must always be copied into that directory on the DOS partition before running /sbin/lilo, e.g. when recompiling a kernel, the standard procedure changes from

# make zlilo

to

# make zImage
# mv /dos/linux/vmlinuz /dos/linux/vmlinuz.old
# mv arch/i386/boot/zImage /dos/linux/vmlinuz
# /sbin/lilo

WARNING: De-fragmenting such a DOS partition is likely to make Linux or even the whole system unbootable. Therefore, the DOS partition should either not be de-fragmented, or a Linux boot disk should be prepared (and tested) to bring up Linux and to run /sbin/lilo after the de-fragmentation.^{11}

11Setting the “system” attribute from DOS on the critical files (e.g. everything in C:\LINUX) may help to protect them from being rearranged. However, the boot floppy should still be ready, just in case.

2 The boot prompt

Immediately after it’s loaded, LILO checks whether one of the following is happening:

• any of the [Shift], [Control] or [Alt] keys is pressed.

• [CapsLock] or [ScrollLock] is set.

If this is the case, LILO displays the boot: prompt and waits for the name of a boot image (i.e. Linux kernel or other operating system). Otherwise, it boots the default boot image^{12} or – if a delay has been specified – waits for one of the listed activities until that amount of time has passed.

At the boot prompt, the name of the image to boot can be entered. Typing errors can be corrected with [BackSpace], [Delete], [Ctrl U] and [Ctrl X]. A list of known images can be obtained by pressing [?] or [Tab].

If [Enter] is pressed and no file name has been entered, the default image is booted.

2.1 Boot command-line options

LILO is also able to pass command-line options to the kernel. Command-line options are words that follow the name of the boot image and that are separated by spaces.

Example:

boot: linux single root=200

This document only gives an overview of boot options. Please consult Paul Gortmaker’s BootPrompt-HOWTO for a more complete and more up to date list. You can get it from

ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO/BootPrompt-HOWTO.gz or from one of the many mirror

sites.

2.1.1 Standard options

Recent kernels recognize a large number of options, among them are debug, no387, no-hlt, ramdisk=size, reserve=base,size,…, root=device, ro, and rw. All current init programs also recognize the option single. The options lock and vga are processed by the boot loader itself. Boot command-line options are always case-sensitive.

single boots the system in single-user mode. This bypasses most system initialization procedures and directly starts a root shell on the console. Multi-user mode can typically be entered by exiting the singleuser shell or by rebooting.

root=device changes the root device. This overrides settings that may have been made in the boot image and on the LILO command line. device is either the hexadecimal device number ^{13} or the full path name of the device, e.g. /dev/hda3.^{14}

reserve=base,size,… reserves IO port regions. This can be used to prevent device drivers from auto-probing addresses where other devices are located, which get confused by the probing. ro instructs the kernel to mount the root file system read-only. rw mounts it read-write. If neither ro nor rw is specified, the setting from the boot image is used.

12The default boot image is either the first boot image, the image specified with the default variable, or the image that has been selected at the boot prompt.

13This is a list of device numbers of some frequently used devices:

/dev/fd0 200 /dev/hda1 301 /dev/sda1 801

/dev/fd1 201 /dev/hda2 302 /dev/sda2 802

⋯ ⋯ ⋯

/dev/hdb1 341 /dev/sdb1 811 /dev/hdb2 342 /dev/sdb2 812

⋯  ⋯ 14The device names are hard-coded in the kernel. Therefore, only the “standard” names are supported and some less common devices may not be recognized. In those cases, only numbers can be used.

no-hlt avoids executing a HLT instructions whenever the system is idle. HLT normally significantly reduces power consumption and therefore also heat dissipation of the CPU, but may not work properly with some clone CPUs. no387 disables using the hardware FPU even if one is present.

debug enables more verbose console logging.

Recent kernels also accept the options init=name and noinitrd. init specifies the name of the init program to execute. Therefore, if single mode cannot be entered because init is mis-configured, one may still be able to reach a shell using init=/bin/sh. noinitrd disables automatic loading of the initial RAM disk. Instead, its content is then available on /dev/initrd.

vga=mode alters the VGA mode set at startup. The values normal, extended, ask or a decimal number are recognized. (See also page 25.)

kbd=code,… preloads a sequence of keystrokes in the BIOS keyboard buffer. The keystrokes have to be entered as 16 bit hexadecimal numbers, with the upper byte containing the scan code and the lower byte containing the ASCII code. Note that most programs only use the ASCII code, so the scan code can frequently be omitted. Scan code tables can be found in many books on PC hardware. Note that scan codes depend on the keyboard layout.

Finally, lock stores the current command-line as the default command-line, so that LILO boots the same image with the same options (including lock) when invoked the next time.

2.1.2 Device-specific options

There is also a plethora of options to specify certain characteristics (e.g. IO and memory addresses) of devices. Some common ones are ether, floppy, hd, bmouse, and sound. The usage of these options is option=number,. . .. Please consult the corresponding FAQs and HOWTOs for details. For an overview of all available options, consult the file init/main.c in the kernel source tree.

2.1.3 Other options

Options of the type variable=value which are neither standard options nor device-specific options, cause the respective variables to be set in the environment passed to init. The case of the variable name is preserved, i.e. it isn’t automatically converted to upper case.

Note that environment variables passed to init are typically available in system initialization scripts (e.g. /etc/rc.local), but they’re not visible from ordinary login sessions, because the login program removes them from the user’s environment.

2.1.4 Repeating options

The effect of repeating boot command-line options depends on the options.^{15} There are three possible behaviours:

Options that only enable or disable a certain functionality can be repeated any number of times. debug, lock, no-hlt, and no387 fall into this category.

Other options change a global setting whenever they appear, so only the value or presence of the last option matters. The antagonists ro and rw are such options. Also, ramdisk, root, and vga work this way. Example: ro rw would mount the root file system read-write.

15Options are frequently repeated when a string defined with append or literal is prepended to the parameters typed in by the user. Also, LILO implicitly prepends the options ramdisk, ro, root, or rw when ramdisk, read-only, read-write, or root, respectively, are set in the configuration file. (lock and vga are handled by a different internal mechanism.)

Finally, when reserve and many device-specific options are repeated, each occurrence has its own meaning, e.g. hd=… hd=… would configure two hard disks, and reserve=0x300,8 reserve=0x5f0,16 would reserve the ranges 0x300 to 0x307 and 0x5f0 to 0x5ff (which is equivalent to writing reserve=0x300,8,0x5f0,16).

2.1.5 Implicit options

LILO always passes the string BOOT IMAGE=name to the kernel, where name is the name by which the kernel is identified (e.g. the label). This variable can be used in /etc/rc to select a different behaviour, depending on the kernel.

When booting automatically, i.e. without human intervention, the word auto is also passed on the command line. This can be used by init to suppress interactive prompts in the boot phase.

2.2 Boot image selection

The details of selecting the boot image are somewhat complicated. The following tables illustrate them. First, if neither prompt is set nor a shift key is being pressed:

Externally Command Auto- Booted image
provided line in matic
cmd. line^{16} map file^{17} boot^{18}
No  No  Yes Default image
Yes  —  Yes Specified by external

command line

No Yes Yes Specified by command line

in map file

If prompt is not set and a shift key is being pressed:

Input Empty Extern. Cmd.l. Auto- Booted image
timeout cmd.l. cmd.l. in map matic
file  boot
No  No  —  —  No Specified by the user
No  Yes  —  —  No Default image
Yes  n/a  —  —  Yes Default image

Finally, if the configuration variable prompt is set:

Input Empty Extern. Cmd.l. Auto- Booted image
timeout cmd.l. cmd.l. in map matic
file  boot
No  No  No  No  No Specified by the user
No  Yes  No  No  No Default image
Yes  n/a  No  No  Yes Default image
n/a  n/a  Yes  —  Yes Specified by external

command line

n/a n/a No Yes Yes Specified by command

line in map file Note that LILO pauses for the amount of time specified in delay when at the end of a default command line. The automatic boot can then be interrupted by pressing a modifier key ([Shift], [Ctrl], etc.).

16Externally provided command lines could be used to add front-ends to LILO. They would pass the respective command string to LILO, which would then interpret it like keyboard input. This feature is currently not used.

17This command line is set by invoking the map installer with the -R option, by using the boot command-line option lock, or if a fallback command line is set (with fallback).

18I.e. the keyword auto is added.

The default image is the first image in the map file or the image specified with the default variable. However, after an unsuccessful boot attempt, the respective image becomes the default image.

3 Map installer

The map installer program /sbin/lilo updates the boot sector and creates the map file. If the map installer detects an error, it terminates immediately and does not touch the boot sector and the map file.

Whenever the map installer updates a boot sector, the original boot sector is copied to /boot/boot.number, where number is the hexadecimal device number. If such a file already exists, no backup copy is made. Similarly, a file /boot/part.number is created if LILO modifies the partition table. (See page 22.)

3.1 Command-line options

The LILO map installer can be invoked in the following ways:

3.1.1 Show current installation

The currently mapped files are listed. With -v, also many parameters are shown.

/sbin/lilo [ -C config file ] -q [ -m map file ] [ -v …]

-C config file Specifies the configuration file that is used by the map installer (see section 3.2). If -C is omitted,

     /etc/lilo.conf is used.

-m map file Specifies an alternate map file. See also sections 3.1.7 and 3.2.2.

-q

Lists the currently mapped files.

-v … Increase verbosity. See also sections 3.1.7 and 3.2.2.

3.1.2 Create or update map

A new map is created for the images described in the configuration file /etc/lilo.conf and they are registered in the boot sector.

/sbin/lilo [ -C config file ] [ -b boot device ] [ -c ] [ -l ] [ -i boot sector ] [ -f disk tab ] [ -m map file ]

[ -d delay ] [ -v …] [ -t ] [ -s save file | -S save file ] [ -P fix | -P ignore ] [ -r root dir ]

-b boot device Specifies the boot device. See also sections 3.1.7 and 3.2.2.

-c

Enables map compaction. See also sections 3.1.7 and 3.2.2.

-C config file Specifies an alternate configuration file. See also section 3.1.1.

-d delay Sets the delay before LILO boots the default image. Note that the delay is specified in tenths of a second. See also sections 3.1.7 and 3.2.2.

-D name Specifies the default image. See also sections 3.1.7 and 3.2.2.

-f disk tab Specifies a disk parameter table file. See also sections 3.1.7 and 3.2.2.

-i boot sector Specifies an alternate boot file. See also sections 3.1.7 and 3.2.2.

-l

Enables linear sector addresses. See also sections 3.1.7 and 3.2.2.

-m map file Specifies an alternate map file. See also sections 3.1.7 and 3.2.2.

-P mode Specifies how invalid partition table entries should be handled. See also sections 3.1.7 and 3.2.2.

-r root directory Chroots to the specified directory before doing anything else. This is useful when running the map installer while the normal root file system is mounted somewhere else, e.g. when recovering from an installation failure with a recovery disk. The -r option is implied if the environment variable ROOT is set.^{19} The current directory is changed to the new root directory, so using relative paths may not work.

-s save file Specifies an alternate boot sector save file. See also sections 3.1.7 and 3.2.2.

-S save file Like -s, but overwrites old save files.

-t

Test only. This performs the entire installation procedure except replacing the map file, writing the modified boot sector and fixing partition tables. This can be used in conjunction with the -v option to verify that LILO will use sane values.

-v … Increase verbosity. See also sections 3.1.7 and 3.2.2.

3.1.3 Change default command line

Changes LILO’s default command line. See also section 2.2.

/sbin/lilo [ -C config file ] [ -m map file ] -R [ word …]

-C config file Specifies an alternate configuration file. See also section 3.1.1.

-m map file Specifies an alternate map file. See also sections 3.1.7 and 3.2.2.

-R word … Stores the specified words in the map file. The boot loader uses those words as the default command line when booting the next time. That command line is removed from the map file by the boot loader by overwriting the sector immediately after reading it. The first word has to be the name of a boot image. If -R is not followed by any words, the current default command line in the map file is erased.^{20} If the command line isn’t valid, the map installer issues an error message and returns a non-zero exit code.

19E.g. if your root partition is mounted on /mnt, you can update the map by simply running ROOT=/mnt /mnt/sbin/lilo

3.1.4 Kernel name translation

Determines the path of the kernel.

/sbin/lilo [ -C config file ] -I name [ options ]

-C config file Specifies an alternate configuration file. See also section 3.1.1.

-I name [ options ] Translates the specified label name to the path of the corresponding kernel image and prints that path on standard output. This can be used to synchronize files that depend on the kernel (e.g. the ps database). The image name can be obtained from the environment variable BOOT IMAGE. An error message is issued and a non-zero exit code is returned if no matching label name can be found. The existence of the image file is verified if the option character v is added.

3.1.5 De-installation

Restores the boot sector that was used before the installation of LILO. Note that this option only works properly if LILO’s directories (e.g. /boot) have not been touched since the first installation. See also section 4.3.

/sbin/lilo [ -C config file ] [ -s save file ] -u | -U [ boot device ]

-C config file Specifies an alternate configuration file. See also section 3.1.1.

-s save file Specifies an alternate boot sector save file. See also sections 3.1.7 and 3.2.2.

-u [ device name ] Restores the backup copy of the specified boot sector. If no device is specified, the value of the boot variable is used. If this one is also unavailable, LILO uses the current root device. The name of the backup copy is derived from the device name. The -s option or the backup variable can be used to override this. LILO validates the backup copy by checking a time stamp.

-U [ device name ] Like -u, but does not check the time stamp.

3.1.6 Print version number

/sbin/lilo -V
-V

Print the version number and exit.

20-R is typically used in reboot scripts, e.g.

#!/bin/sh cd / /sbin/lilo -R "$*" && reboot

3.1.7 Options corresponding to configuration variables

There are also many command-line options that correspond to configuration variables. See section 3.2.2 for a description.

Command-line option Configuration variable
-b boot device  boot=boot device

-c                  compact

-d tsecs  delay=tsecs

-D name  default=name

-i boot sector  install=boot sector

-l                  linear

-m map file  map=map file

-P fix              fix-table
-P ignore           ignore-table

-s backup file  backup=backup file
-S backup file  force-backup=backup file

-v . . .  verbose=level

3.2 Configuration

The configuration information is stored in the file /etc/lilo.conf and consists of variable assignments.

3.2.1 Syntax

The following syntax rules apply:

• flag variables consist of a single word and are followed by whitespace or the end of the file.

• string variables consist of the variable name, optional whitespace, an equal sign, optional whitespace, the value and required whitespace, or the end of the file.

• a non-empty sequence of blanks, tabs, newlines and comments counts as whitespace.

• variable names are case-insensitive. Values are usually case-sensitive, but there are a few exceptions. (See below.)

• tabs and newlines are special characters and may not be part of a variable name or a value. The use of other control characters and non-ASCII characters is discouraged.

• blanks and equal signs may only be part of a variable name or a value if they are escaped by a backslash or if the value is embedded in double quotes. An equal sign may not be the only character in a name or value.

• an escaped tab is converted to an escaped blank. An escaped newline is removed from the input stream. An escaped backslash (i.e. two backslashes) is converted to a backslash. Inside quoted strings, only double quotes, backslashes, dollar signs, and newlines can be escaped.

• quoted strings can be continued over several lines by ending each incomplete line with a backslash. A single space is inserted in the string for the line end and all spaces or tabs that follow immediately on the next line.

• environment variables can be used by specifying them in the form $name or ${name}. Dollar signs can be escaped.

• comments begin with a number sign and end with the next newline. All characters (including backslashes) until the newline are ignored.

Example:

boot = $FD
install = $MNT/boot.b
map = $MNT/map
compact
read-only
append = "nfsroot=/home/linux-install/root \
  nfsaddrs=128.178.156.28:128.178.156.24::255.255.255.0:lrcinst"
image = $MNT/zImage

3.2.2 Global options

/etc/lilo.conf begins with a possibly empty global options section. Many global options can also be set from the command line, but storing permanent options in the configuration file is more convenient.

The following global options are recognized:

backup=backup file Copy the original boot sector to backup file (which may also be a device, e.g. /dev/null) instead of /boot/boot.number

boot=boot device Sets the name of the device (e.g. a hard disk partition) that contains the boot sector. If boot is omitted, the boot sector is read from (and possibly written to) the device that is currently mounted as root.

change-rules Defines partition type numbers. See section 3.5.2 for details.

compact Tries to merge read requests for adjacent sectors into a single read request. This drastically reduces load time and keeps the map smaller. Using compact is especially recommended when booting from a floppy disk. compact may conflict with linear, see section 5.3.

default=name Uses the specified image as the default boot image. If default is omitted, the image appearing first in the configuration file is used.

delay=tsecs Specifies the number of tenths of a second LILO should wait before booting the first image. This is useful on systems that immediately boot from the hard disk after enabling the keyboard. LILO doesn’t wait if delay is omitted or if delay is set to zero.

disk=device name Defines non-standard parameters for the specified disk. See section 3.4 for details.

fix-table Allows LILO to adjust 3D addresses in partition tables. Each partition entry contains a 3D (sector/head/cylinder) and a linear address of the first and the last sector of the partition. If a partition is not track-aligned and if certain other operating systems (e.g. PC/MS-DOS or OS/2) are using the same disk, they may change the 3D address. LILO can store its boot sector only on partitions where both address types correspond. LILO re-adjusts incorrect 3D start addresses if fix-table is set.

WARNING: This does not guarantee that other operating systems may not attempt to reset the address later. It is also possible that this change has other, unexpected side-effects. The correct fix is to re-partition the drive with a program that does align partitions to tracks. Also, with some disks (e.g. some large EIDE disks with address translation enabled), under some circumstances, it may even be unavoidable to have conflicting partition table entries.

force-backup=backup file Like backup, but overwrite an old backup copy if it exists. backup=backup file is ignored if force-backup appears in the same configuration file. ignore-table Tells LILO to ignore corrupt partition tables and to put the boot sector even on partitions

that appear to be unsuitable for that.

install=boot sector Install the specified file as the new boot sector. If install is omitted, /boot/boot.b is used as the default.

keytable=table file Re-map the keyboard as specified in this file. See section 3.6 for details.

linear Generate linear sector addresses instead of sector/head/cylinder addresses. Linear addresses are translated at run time and do not depend on disk geometry. Note that boot disks may not be portable if linear is used, because the BIOS service to determine the disk geometry does not work reliably for floppy disks. When using linear with large disks, /sbin/lilo may generate references to inaccessible disk areas (see section 1.3.1), because 3D sector addresses are not known before boot time. linear may conflict with compact, see section 5.3.

map=map file Specifies the location of the map file. If map is omitted, a file /boot/map is used.

message=message file Specifies a file containing a message that is displayed before the boot prompt. No message is displayed while waiting for a modifier key ([Shift], etc.) after printing “LILO ”. In the message, the FF character ([Ctrl L]) clears the local screen. The size of the message file is limited to 65535 bytes. The map file has to be rebuilt if the message file is changed or moved.

nowarn Disables warnings about possible future dangers.

prompt Forces entering the boot prompt without expecting any prior key-presses. Unattended reboots are impossible if prompt is set and timeout isn’t.

serial=parameters Enables control from a serial line. The specified serial port is initialized and LILO is accepting input from it and from the PC’s keyboard. Sending a break on the serial line corresponds to pressing a shift key on the console in order to get LILO’s attention. All boot images should be password-protected if the serial access is less secure than access to the console, e.g. if the line is connected to a modem. The parameter string has the following syntax: port,bps parity bits The components bps, parity and bits can be omitted. If a component is omitted, all following components have to be omitted too. Additionally, the comma has to be omitted if only the port number is specified.

port the number of the serial port, zero-based. 0 corresponds to COM1 alias /dev/ttyS0, etc. All four

ports can be used (if present). bps the baud rate of the serial port. The following baud rates are supported: 110, 300, 1200, 2400,

4800, 9600, 19200, and 38400 bps. Default is 2400 bps. parity the parity used on the serial line. LILO ignores input parity and strips the 8th bit. The following

(upper or lower case) characters are used to describe the parity: n for no parity, e for even parity and o for odd parity.

bits the number of bits in a character. Only 7 and 8 bits are supported. Default is 8 if parity is “none”, 7 if parity is “even” or “odd”.

If serial is set, the value of delay is automatically raised to 20.

Example: serial=0,2400n8 initializes COM1 with the default parameters.

timeout=tsecs Sets a timeout (in tenths of a second) for keyboard input. If no key is pressed for the specified time, the first image is automatically booted. Similarly, password input is aborted if the user is idle for too long. The default timeout is infinite.

verbose=level Turns on lots of progress reporting. Higher numbers give more verbose output. If -v is additionally specified on the command line, level is increased accordingly. The following verbosity levels exist:

<0 only warnings and errors are shown

0 prints one line for each added or skipped image

1 mentions names of important files and devices and why they are accessed. Also displays informational messages for exceptional but harmless conditions and prints the version number.

2 displays statistics and processing of temporary files and devices

3 displays disk geometry information and partition table change rules

4 lists sector mappings as they are written into the map file (i.e. after compaction, in a format suitable to pass it to the BIOS)

5 lists the mapping of each sector (i.e. before compaction, raw)

When using the -q option, the levels have a slightly different meaning:

0 displays only image names

1 also displays all global and per-image settings

2 displays the address of the first map sector

Additionally, the kernel configuration parameters append, initrd, ramdisk, read-only, read-write, root

and vga, and the general per-image options fallback, lock, optional, password, restricted, and

single-key can be set in the global options section. They are used as defaults if they aren’t specified in the configuration sections of the respective images. See below for a description.

The plethora of options may be intimidating at first, but in “normal” configurations, hardly any options but

boot, compact, delay, root, and vga are used.

3.2.3 General per-image options

The following options can be specified for all images, independent of their type:

alias=name Specifies a second name for the current entry.

fallback=command line Specifies a string that is stored as the default command line if the current image is booted. This is useful when experimenting with kernels which may crash before allowing interaction with the system. If using the fallback option, the next reboot (e.g. triggered by a manual reset or by a watchdog timer) will load a different (supposedly stable) kernel. The command line by the fallback mechanism is cleared by removing or changing the default command line with the -R option, see 3.1.3.

label=name By default, LILO uses the main file name (without its path) of each image specification to identify that image. A different name can be used by setting the variable label.

lock Enables automatic recording of boot command lines as the defaults for the following boots. This way, LILO “locks” on a choice until it is manually overridden.

optional Omit this image if its main file is not available at map creation time. This is useful to specify test kernels that are not always present.

password=password Ask the user for a password when trying to load this image. Because the configuration file contains unencrypted passwords when using this option, it should only be readable for the superuser. Passwords are always case-sensitive.

restricted Relaxes the password protection by requiring a password only if parameters are specified on the command line (e.g. single). restricted can only be used together with password. single-key Enables booting the image by hitting a single key, without the need to press [Enter] afterwards.

single-key requires that either the image’s label or its alias (or both) is a single character. Furthermore, no other image label or alias may start with that character, e.g. an entry specifying a label linux and an alias l is not allowed with single-key. Note that you can’t specify command-line parameters for an entry for which only single-keyed names exist.

All general per-image options, with the exception of label and alias, can also be set in the global options section as defaults for all images.

Example:

password = Geheim
single-key
image = /vmlinuz
  label = linux
  alias = 1
  restricted
other = /dev/hda1
  label = dos
  alias = 2

3.2.4 Per-image options for kernels

Each (kernel or non-kernel) image description begins with a special variable (see section 3.3) which is followed by optional variables. The following variables can be used for all image descriptions that describe a Linux kernel:

append=string Appends the options specified in string to the parameter line passed to the kernel. This is typically used to specify parameters of hardware that can’t be entirely auto-detected, e.g.

     append = "hd=64,32,202"

initrd=name Specifies the file that will be loaded at boot time as the initial RAM disk.

literal=string like append, but removes all other options (e.g. setting of the root device). Because vital options can be removed unintentionally with literal, this option cannot be set in the global options section.

ramdisk=size Specifies the size of the optional RAM disk. A value of zero indicates that no RAM disk should be created. If this variable is omitted, the RAM disk size configured into the boot image is used.

read-only Specifies that the root file system should be mounted read-only. Typically, the system startup procedure re-mounts the root file system read-write later (e.g. after fsck’ing it).

read-write specifies that the root file system should be mounted read-write.

root=root device Specifies the device that should be mounted as root. If the special name current is used, the root device is set to the device on which the root file system is currently mounted. If the root has been changed with -r, the respective device is used. If the variable root is omitted, the root device setting contained in the kernel image is used. It can be changed with the rdev program.

vga=mode Specifies the VGA text mode that should be selected when booting. The following values are recognized (case is ignored):

normal select normal 80x25 text mode.

extended select 80x50 text mode. The word extended can be abbreviated to ext. ask stop and ask for user input (at boot time). number use the corresponding text mode. A list of available modes can be obtained by booting with

vga=ask and pressing [Enter].

If this variable is omitted, the VGA mode setting contained in the kernel image is used. rdev supports manipulation of the VGA text mode setting in the kernel image.

All kernel per-image options but literal can also be set in the global options section as defaults for all kernels.

If one of ramdisk, read-only, read-write, root, or vga is omitted in the configuration file and the corresponding value in the kernel image is changed, LILO or the kernel will use the new value.

It is perfectly valid to use different settings for the same image, because LILO stores them in the image descriptors and not in the images themselves.

Example:

image = /vmlinuz
  label = lin-hd
  root = /dev/hda2
image = /vmlinuz
  label = lin-fd
  root = /dev/fd0

3.3 Boot image types

LILO can boot the following types of images:

• kernel images from a file.

• kernel images from a block device. (E.g. a floppy disk.)

• the boot sector of some other operating system.

The image type is determined by the name of the initial variable of the configuration section.

The image files can reside on any media that is accessible at boot time. There’s no need to put them on the root device, although this certainly doesn’t hurt.

3.3.1 Booting kernel images from a file

The image is specified as follows: image=name

Example:

image = /linux

See sections 3.2.3 and 3.2.4 for the options that can be added in a kernel image section.

3.3.2 Booting kernel images from a device

The range of sectors that should be mapped has to be specified. Either a range (start-end) or a start and a distance (start+number) have to be specified. start and end are zero-based. If only the start is specified, only that sector is mapped.

The image is specified as follows: image=device name Additionally, the range variable must be set.

Example:

image = /dev/fd0
  range = 1+512

All kernel options can also be used when booting the kernel from a device.

3.3.3 Booting a foreign operating system

LILO can even boot other operating systems, i.e. MS-DOS. To boot an other operating system, the name of a loader program, the device or file that contains the boot sector and the device that contains the partition table have to be specified.

The boot sector is merged with the partition table and stored in the map file.

Currently, the loaders chain.b and os2 d.b exist. chain.b simply starts the specified boot sector.^{21} os2 d.b it a variant of chain.b that can boot OS/2 from the second hard disk. The map-drive option has to be used with os2 d.b to actually swap the drives.

The image is specified as follows: other=device name or other=file name

In addition to the options listen in section 3.2.3, the following variables are recognized:

change Change the partition table according to the rules specified in this change section. This option is intended for booting systems which find their partitions by examining the partition table. See section 3.5.3 for details.

loader=chain loader Specifies the chain loader that should be used. If it is omitted, /boot/chain.b is used.

map-drive=bios device code Instructs chain.b to installs a resident driver that re-maps the floppy or hard disk drives. This way, one can boot any operating system from a hard disk different from the first one, as long as that operating system uses only the BIOS to access that hard disk.^{22} This is known to work for PC/MS-DOS.

map-drive is followed by the variable to=bios device code which specifies the drive that should effectively be accessed instead of the original one. The list of mappings is only searched until the first match is found. It is therefore possible to “swap” drives, see the second example below.

table=device Specifies the device that contains the partition table. LILO does not pass partition information to the booted operating system if this variable is omitted. (Some operating systems have other means to determine from which partition they have been booted. E.g. MS-DOS usually stores the geometry of the boot disk or partition in its boot sector.) Note that /sbin/lilo must be re-run if a partition table mapped referenced with table is modified.

unsafe Do not access the boot sector at map creation time. This disables some sanity checks, including a partition table check. If the boot sector is on a fixed-format floppy disk device, using unsafe avoids the need to put a readable disk into the drive when running the map installer. unsafe and table are mutually incompatible.

None of these options can be set in the global options section.

Examples:

other = /dev/hda2
  label = dos
  table = /dev/hda

21The boot sector is loaded by LILO’s secondary boot loader before control is passed to the code of chain.b.

22So you should be very suspicious if the operating system requires any specific configuration or even drivers to use the disk it is booted from. Since there is a general trend to use optimized drivers to fully exploit the hardware capabilities (e.g.

non-blocking disk access), booting systems from the second disk may become increasingly difficult.

other = /dev/hdb2
  label = os2
  loader = /boot/os2_d.b
  map-drive = 0x80
    to = 0x81
  map-drive = 0x81
    to = 0x80

3.4 Disk geometry

For floppies and most hard disks, LILO can obtain the disk geometry information from the kernel. Unfortunately, there are some exotic disks or adapters which may either not supply this information or which may even return incorrect information.

If no geometry information is available, LILO reports either the error

geo query dev HDIO GETGEO (dev 0xnumber)

or

Device 0xnumber:  Got bad geometry sec/hd/cyl

If incorrect information is returned, booting may fail in several ways, typically with a partial “LILO” banner message. In this document, that is called a “geometry mismatch”.

The next step should be to attempt setting the linear configuration variable or the -l command-line option. If this doesn’t help, the entire disk geometry has to be specified explicitly. Note that linear doesn’t always work with floppy disks.

Another common use of disk sections is if an (E)IDE and a SCSI drive are used in the same system and the BIOS is configured to use the SCSI drive as the first drive. (Normally, the (E)IDE drive would be the first drive and the SCSI drive would be the second one.) Since LILO doesn’t know how the BIOS is configured, it needs to be told explicitly about this arrangement. (See the second example below.)

3.4.1 Obtaining the geometry

The disk geometry parameters can be obtained by booting MS-DOS and running the program DPARAM.COM with the hexadecimal BIOS code of the drive as its argument, e.g. dparam 0x80 for the first hard disk. It displays the number of sectors per track, the number of heads per cylinder and the number of cylinders. All three numbers are one-based.

Alternatively, the geometry may also be determined by reading the information presented by the “setup” section of the ROM-BIOS or by using certain disk utilities under operating systems accessing the disk through the BIOS.

3.4.2 Specifying the geometry

Disk geometry parameters are specified in the options section of the configuration file. Each disk parameter section begins with disk=disk device, similar to the way how boot images are specified. It is suggested to group disk parameter sections together, preferably at the beginning or the end of the options section.

For each disk, the following variables can be specified:

bios=bios device code Is the number the BIOS uses to refer to that device. Normally, it’s 0x80 for the first hard disk and 0x81 for the second hard disk. Note that hexadecimal numbers have to begin with “0x”. If bios is omitted, LILO tries to “guess” that number.

sectors=sectors and heads=heads specify the number of sectors per track and the number of heads, i.e. the number of tracks per

cylinder. Both parameters have to be either specified together or they have to be entirely omitted. If omitted, LILO tries to obtain that geometry information from the kernel.

cylinders=cylinders Specifies the number of cylinders. This value is only used for sanity checks. If cylinders is omitted, LILO uses the information obtained from the kernel if geometry information had to be requested in order to determine some other parameter. Otherwise,^{23} it just assumes the number of cylinders to be 1024, which is the cylinder limit imposed by the BIOS.

inaccessible Marks the device as inaccessible (for the BIOS). This is useful if some disks on the system can’t be read by the BIOS, although LILO “thinks” they can. If one accidentally tries to use files located on such disks for booting, the map installer won’t notice and the system becomes unbootable. The most likely use of inaccessible is to prevent repetition after experiencing such a situation once. No other variables may be specified if a device is configured as inaccessible.

Additionally, partition subsections can be added with partition=partition device. Each partition section can contain only one variable:

start=partition offset Specifies the zero-based number of the start sector of that partition. The whole disk always has a partition offset of zero. The partition offset is only necessary when using devices for which the kernel does not provide that information, e.g. CD-ROMs.

Examples:

disk = /dev/sda
  bios = 0x80
  sectors = 32
  heads = 64
  cylinders = 632
  partition = /dev/sda1
    start = 2048
  partition = /dev/sda2
    start = 204800
  partition = /dev/sda3
    start = 500000
  partition = /dev/sda4
    start = 900000
disk = /dev/sda
  bios = 0x80
disk = /dev/hda
  bios = 0x81

3.5 Partition table manipulation

Some non-Linux operating systems obtain information about their partitions (e.g. their equivalent of the root file system) from the partition table. If more than one such operating system is installed on a PC, they may have conflicting interpretations of the content of the partition table. Those problems can be avoided by changing the partition table, depending on which operating system is being booted.

Partition table changes are specified in a change section in the configuration file section describing the foreign operating system. Note that change sections are only accepted if the build-time option REWRITE TABLE is set.

23I.e. if the BIOS device code, the number of sectors, the number of heads and the partition start are specified. Note that the number of cylinders may appear to vary if cylinders is absent and only some of the partition starts are specified.

The change section contains subsections for each partition whose table entry needs to be modified. Partitions are specified with partition=device name

Changes are applied in the sequence in which they appear in the configuration file. Configurations containing changes that are redundant either by repeating a previous change or by changing its result further are invalid and refused by the map installer.

Internally, all changes are expressed as rules which specify the location (disk and offset in the partition table), the value this location must contain before the change, and the value that has to be stored. As a safety measure, the rule is ignored if the previous value is found to be different.

3.5.1 Partition activation

This option is intended for booting systems which determine their boot partition by examining the active flag in the partition table. The flag is enabled with activate and disabled with deactivate. Note that only the current partition is affected. LILO does not automatically change the active flags of other partitions and it also allows more than one partition to be active at the same time.

Example:

other = /dev/sda4
  label = sco
  change
    partition = /dev/sda4
      activate
    partition = /dev/sda3
      deactivate

3.5.2 Partition type change rules

Partition type changes are normally a transition between two possible values, e.g. a typical convention is to set the highest bit in the partition type in order to “hide”, and to clear it to “unhide” a partition. LILO performs these changes based on a set of rules. Each rule defines the name of a partition type, its normal value, and the value when hidden. Those rules are defined in the options section of the configuration file. The section defining them begins with change-rules.

The following options and variables can appear in the section:

reset Removes all previously defined rules. This is needed if a user doesn’t wish to use the pre-defined rules (see below).

type=name Adds a rule for the type with the specified name. Type names are case-insensitive. The values are defined with normal=byte and hidden=byte. Values can be specified as decimal or as hexadecimal numbers with a leading 0x. If only one of the values is present, the other value is assumed to be the same number, but with the most significant bit inverted.

LILO pre-defines rules for the three partition types of DOS partitions. The following example removes the pre-defined rules and creates them again:

change-rules
  reset
  type = DOS12
    normal = 0x01
    hidden = 0x81
  type = DOS16_small
    normal = 4 # hidden is 0x84

  type = DOS16_big
    hidden = 0x86

3.5.3 Partition type changes

Partition type changes are specified in the partition section as set=name state, where name is the name of the partition type, and state is its state, i.e. normal or hidden.

Example:

other = /dev/sda3
  label = dos
  change
    partition = /dev/sda2
      set = dos16_big_normal
    partition = /dev/sda3
      activate
      set = DOS16_big_normal

Only one set variable is allowed per partition section. In the rare event that more than one set variable is needed, further partition sections can be used.

3.6 Keyboard translation

The PC keyboard emits so-called scan codes, which are basically key numbers. The BIOS then translates those scan codes to the character codes of the characters printed on the key-caps. By default, the BIOS normally assumes that the keyboard has a US layout. Once an operating system is loaded, this operating system can use a different mapping.

At boot time, LILO only has access to the basic services provided by the BIOS and therefore receives the character codes for an US keyboard. It provides a simple mechanism to re-map the character codes to what is appropriate for the actual layout.^{24}

3.6.1 Compiling keyboard translation tables

LILO obtains layout information from the keyboard translation tables Linux uses for the text console. They are usually stored in /usr/lib/kbd/keytables. LILO comes with a program keytab-lilo.pl that reads those tables and generates a table suitable for use by the map installer. keytab-lilo.pl invokes the program loadkeys to print the tables in a format that is easy to parse.^{25}

keytab-lilo.pl is used as follows:

keytab-lilo.pl [ -p old code=new code ] … [path]default layout[.extension] ] [path]kbd layout[.extension] ]

-p old code=new code Specifies corrections (“patches”) to the mapping obtained from the translation table files. E.g. if pressing the upper case “A” should yield an at sign, -p 65=64 would be used. The -p option can be repeated any number of times. The codes can also be given as hexadecimal or as octal numbers if they are prefixed with 0x or 0, respectively. path The directory in which the file resides. The default path is /usr/lib/kbd/keytables. extension Usually the trailing .map, which is automatically added if the file name doesn’t contain dots. default layout Is the layout which specifies the translation by the BIOS. If none is specified, us is assumed. kbd layout Is the actual layout of the keyboard. keytab-lilo.pl writes the resulting translation table as a binary string to standard output. Such tables can be stored anywhere with any name, but the suggested naming convention is /boot/kbd.ktl (“\underline{K}eyboard \underline{T}able for \underline{L}ilo”), where kbd is the name of the keyboard layout. Example:

24The current mechanism isn’t perfect, because it sits on top of the scan code to character code translation performed by the BIOS. This means that key combinations that don’t produce any useful character on the US keyboard will be ignored by LILO. The advantage of this approach is its simplicity.

25On some systems, only root can execute loadkeys. It is then necessary to run keytab-lilo.pl as root too.

keytab-lilo.pl de >/boot/de.ktl

3.6.2 Using keyboard translation tables The keyboard translation table file is specified with the global configuration option keytable=table file. The complete name of the file has to be given. Example:

keytable = /boot/de.klt

4 Installation and updates

4.1 Installation

This section describes the installation of LILO. See section 4.3 for how to uninstall LILO.

4.1.1 Compatibility

The kernel header files have to be in /usr/include/linux and the kernel usually has to be configured by running make config before LILO can be compiled.

/bin/sh has to be a real Bourne shell. bash is sufficiently compatible, but some ksh clones may cause problems.

A file named INCOMPAT is included in the distribution. It describes incompatibilities to older versions of LILO and may also contain further compatibility notes.

4.1.2 Quick installation

If you want to install LILO on your hard disk and if you don’t want to use all its features, you can use the quick installation script. Read QuickInst for details.

QuickInst can only be used for first-time installations or to entirely replace an existing installation, not to update or modify an existing installation of LILO. Be sure you’ve extracted LILO into a directory that doesn’t contain any files of other LILO installations.

4.1.3 Files

Some of the files contained in lilo-20.tar.gz:

lilo/README

This documentation in plain ASCII format. Some sections containing complex tables are only included in the L^{A}TEX version in doc/user.tex

lilo/INCOMPAT

List of incompatibilities to previous versions of LILO.

lilo/CHANGES

Change history.

lilo/VERSION

The version number of the respective release.

lilo/QuickInst

Quick installation script.

lilo/lilo-version.lsm

The LSM (“\underline{L}inux \underline{S}oftware \underline{M}ap”) entry of the respective LILO release.

lilo/Makefile

Makefile to generate everything else.

lilo/*.c, lilo/*.h

LILO map installer C source and common header files.

lilo/*.S

LILO boot loader assembler source.

lilo/activate.c

C source of a simple boot partition setter.

lilo/dparam.s

Assembler source of a disk parameter dumper.

lilo/mkdist

Shell script used to create the current LILO distribution.

lilo/keytab-lilo.pl

Perl script to generate keyboard translation tables.

lilo/doc/README

Description of how to generate the documentation.

lilo/doc/Makefile

Makefile used to convert the L^{A}TEX source into either DVI output or the plain ASCII README file.

lilo/doc/user.tex

L^{A}TEX source of LILO’s user’s guide (this document).

lilo/doc/tech.tex

L^{A}TEX source of LILO’s technical overview.

lilo/doc/*.fig

Various xfig pictures used in the technical overview.

lilo/doc/fullpage.sty

Style file to save a few square miles of forest.

lilo/doc/rlatex

Shell script that invokes L^{A}TEX repeatedly until all references have settled.

lilo/doc/t2a.pl

Perl script to convert the L^{A}TEX source of the user’s guide to plain ASCII.

Files created after make in lilo/ (among others):

lilo/boot.b

Combined boot sector. make install puts this file into /boot

lilo/chain.b

Generic chain loader. make install puts this file into /boot

lilo/os2 d.b

Chain loader to load OS/2 from the second hard disk. make install puts this file into /boot

lilo/lilo

LILO (map) installer. make install puts this file into /sbin

lilo/activate

Simple boot partition setter.

lilo/dparam.com

MS-DOS executable of the disk parameter dumper.

4.1.4 Normal first-time installation

First, you have to install the LILO files:

• extract all files from lilo-version.tar.gz in a new directory.^{26}

• configure the Makefile (see section 4.1.5)

• run make to compile and assemble all parts.

• run make install to copy all LILO files to the directories where they’re installed. /sbin should now contain the file lilo, /usr/sbin should contain keytab-lilo.pl, and /boot should contain

boot.b, chain.b, and os2 d.b.

If you want to use LILO on a non-standard disk, you might have to determine the parameters of your disk(s) and specify them in the configuration file. See section 3.4 for details. If you’re using such a non-standard system, the next step is to test LILO with the boot sector on a floppy disk:

• insert a blank (but low-level formatted) floppy disk into /dev/fd0.

• run echo image=kernel image |

/sbin/lilo -C - -b /dev/fd0 -v -v -v

If you’ve already installed LILO on your system, you might not want to overwrite your old map file. Use the -m option to specify an alternate map file name.

• reboot. LILO should now load its boot loaders from the floppy disk and then continue loading the kernel from the hard disk.

Now, you have to decide, which boot concept you want to use. Let’s assume you have a Linux partition on /dev/hda2 and you want to install your LILO boot sector there. The DOS-MBR loads the LILO boot sector.

• get a working boot disk, e.g. an install or recovery disk. Verify that you can boot with this setup and that you can mount your Linux partition(s) with it.

• if the boot sector you want to overwrite with LILO is of any value (e.g. it’s the MBR or if it contains a boot loader you might want to use if you encounter problems with LILO), you should mount your boot disk and make a backup copy of your boot sector to a file on that floppy, e.g.

dd if=/dev/hda of=/fd/boot_sector bs=512 count=1

• create the configuration file /etc/lilo.conf, e.g.

global settings image specification

per-image options … Be sure to use absolute paths for all files. Relative paths may cause unexpected behaviour when using the -r option.

• now, you can check what LILO would do if you were about to install it on your hard disk:

/sbin/lilo -v -v -v -t

• if you need some additional boot utility (i.e. BOOTACTV), you should install that now

• run /sbin/lilo to install LILO on your hard disk

• if you have to change the active partition, use fdisk or activate to do that

• reboot 4.1.5 Build-time configuration

26E.g. /usr/src/lilo

Certain build-time parameters can be configured. They can either be edited in the top-level Makefile or they can be stored in a file /etc/lilo.defines. Settings in the Makefile are ignored if that file exists.

The following items can be configured:

BEEP Enables beeping after displaying “LILO”. This is useful on machines without speaker and when working over a serial console. This option is disabled by default.

IGNORECASE Makes image name matching case-insensitive, i.e. “linux” and “Linux” are identical. This option is enabled by default. Note that password matching is always case-sensitive.

NO1STDIAG Do not generate diagnostics on read errors in the first stage boot loader. This avoids possibly irritating error codes if the disk controller has transient read problems. This option is disabled by default.

NODRAIN The boot loader empties the keyboard buffer when starting, because it may contain garbage on some systems. Draining the keyboard buffer can be disabled by setting the NODRAIN option. NODRAIN is disabled by default.

NOINSTDEF If the option install is omitted, don’t install a new boot sector, but try to modify the old boot sector instead. This option is disabled by default.

ONE SHOT Disables the command-line timeout (configuration variable timeout) if any key is pressed. This way, very short timeouts can be used if prompt is set. ONE SHOT is disabled by default.

READONLY Disallows overwriting the default command line sector of the map file. This way, command lines set with -R stay in effect until they are explicitly removed. READONLY also disables lock, fallback, and everything enabled by REWRITE TABLE. This option is disabled by default.

REWRITE TABLE Enables rewriting the partition table at boot time. This may be necessary to boot certain operating systems who expect the active flag to be set on their partition or who need changes in partition types. See also section 3.5. This option is dangerous and it is disabled by default.

VARSETUP Enables the use of variable-size setup segments. This option is enabled by default and is only provided to fall back to fixed-size setup segments in the unlikely case of problems when using prehistoric kernels.

XL SECS=sectors Enable support for extra large (non-standard) floppy disks. The number of sectors is set in the BIOS disk parameter table to the specified value. Note that this hack may yield incorrect behaviour on some systems. This option is disabled by default.

/etc/lilo.defines should be used if one wishes to make permanent configuration changes. The usual installation procedures don’t touch that file. Example:

-DIGNORECASE -DONE_SHOT

After changing the build-time configuration, LILO has to be recompiled with the following commands:

make spotless
make

4.1.6 Floppy disk installation

In some cases^{27}, it may be desirable to install LILO on a floppy disk in a way that it can boot a kernel without accessing the hard disk.

The basic procedure is quite straightforward (see also section 1.3.1):

• a file system has to be created on the file system

• the kernel and boot.b have to be copied to the floppy disk

• /sbin/lilo has to be run to create the map file

This can be as easy as

/sbin/mke2fs /dev/fd0
[ -d /fd ] || mkdir /fd
mount /dev/fd0 /fd
cp /boot/boot.b /fd
cp /zImage /fd
echo image=/fd/zImage label=linux |
  /sbin/lilo -C - -b /dev/fd0 -i /fd/boot.b -c -m /fd/map
umount /fd

The command line of /sbin/lilo is a little tricky. -C - takes the configuration from standard input (naturally, one could also write the configuration into a file), -b /dev/fd0 specifies that the boot sector is written to the floppy disk, -i /fd/boot.b takes the first and second stage loaders from the floppy, -c speeds up the load process, and -m /fd/map puts the map file on the floppy too.

4.2 Updates

LILO is affected by updates of kernels, the whole system and (trivially) of LILO itself. Typically, only /sbin/lilo has to be run after any of those updates and everything will be well again (at least as far as LILO is concerned).

4.2.1 LILO update

Before updating to a new version of LILO, you should read at least the file INCOMPAT which describes incompatibilities with previous releases.

After that, the initial steps are the same as for a first time installation: extract all files, configure the Makefile, run make to build the executables and run make install to install the files.

The old versions of boot.b, chain.b, etc. are automatically renamed to boot.old, chain.old, etc. This is done to ensure that you can boot even if the installation procedure does not finish. boot.old, chain.old, etc. can be deleted after the map file is rebuilt.

Because the locations of boot.b, chain.b, etc. have changed and because the map file format may be different too, you have to update the boot sector and the map file. Run /sbin/lilo to do this.

4.2.2 Kernel update

Whenever any of the kernel files that are accessed by LILO is moved or overwritten, the map has to be re-built.^{28} Run /sbin/lilo to do this.

27E.g. if no hard disk is accessible through the BIOS.

28It is advisable to keep a second, stable, kernel image that can be booted if you forget to update the map after a change to your usual kernel image.

The kernel has a make target “zlilo” that copies the kernel to /vmlinuz and runs /sbin/lilo.

4.2.3 System upgrade

Normally, system upgrades (i.e. installation or removal of packages, possibly replacement of a large part of the installed binaries) do not affect LILO. Of course, if a new kernel is installed in the process, the normal kernel update procedure has to be followed (see section 4.2.2). Also, if kernels are removed or added, it may be necessary to update the configuration file.

If LILO is updated by this system upgrade, /sbin/lilo should be run before booting the upgraded system. It is generally a good idea not to rely on the upgrade procedure to perform this essential step automatically.

However, system upgrades which involve removal and re-creation of entire partitions (e.g. /, /usr, etc.) are different. First, they should be avoided, because they bear a high risk of losing other critical files, e.g. the /etc/XF86Config you’ve spent the last week fiddling with. If an upgrade really has to be performed in such a brute-force way, this is equal with total removal of LILO, followed by a new installation. Therefore, the procedures described in the sections 4.3 and 4.2.1 have to be performed. If you’ve forgotten to make a backup copy of /etc/lilo.conf before the destructive upgrade, you might also have to go through section 4.1.4 again.

4.3 LILO de-installation

In order to stop LILO from being invoked when the system boots, its boot sector has to be either removed or disabled. All other files belonging to LILO can be deleted after removing the boot sector, if desired.^{29}

Again, when removing Linux, LILO must be de-installed before (!) its files (/boot, etc.) are deleted. This is especially important if LILO is operating as the MBR.

LILO 14 (and newer) can be de-installed with lilo -u. If LILO 14 or newer is currently installed, but the first version of LILO installed was older than 14, lilo -U may work. When using -U, the warning at the end of this section applies.

If LILO’s boot sector has been installed on a primary partition and is booted by the “standard” MBR or some partition switcher program, it can be disabled by making a different partition active. MS-DOS’ FDISK, Linux fdisk or LILO’s activate can do that.

If LILO’s boot sector is the master boot record (MBR) of a disk, it has to be replaced with a different MBR, typically MS-DOS’ “standard” MBR. When using MS-DOS 5.0 or above, the MS-DOS MBR can be restored with FDISK /MBR. This only alters the boot loader code, not the partition table.

LILO automatically makes backup copies when it overwrites boot sectors. They are named /boot/boot.nnnn, with nnnn corresponding to the device number, e.g. 0300 is /dev/hda, 0800 is /dev/sda, etc. Those backups can be used to restore the old MBR if no easier method is available. The commands are

dd if=/boot/boot.0300 of=/dev/hda bs=446 count=1 or
dd if=/boot/boot.0800 of=/dev/sda bs=446 count=1

respectively.

WARNING: Some Linux distributions install boot.nnnn files from the system where the distribution was created. Using those files may yield unpredictable results. Therefore, the file creation date should be carefully checked.

4.4 Installation of other operating systems

Some other operating systems (e.g. MS-DOS 6.0) appear to modify the MBR in their install procedures. It is therefore possible that LILO will stop to work after such an installation and Linux has to be booted from floppy disk. The original state can be restored by either re-running /sbin/lilo (if LILO is installed as the MBR) or by making LILO’s partition active (if it’s installed on a primary partition).

29Backup copies of old boot sectors may be needed when removing the boot sector. They are stored in /boot.

It is generally a good idea to install LILO after the other operating systems have been installed. E.g. OS/2 is said to cause trouble when attempting to add it to an existing Linux system. (However, booting from floppy and running /sbin/lilo should get around most interferences.)

Typically, the new operating system then has to be added to LILO’s configuration (and /sbin/lilo has to be re-run) in order to boot it.

See also section 5.3 for a list of known problems with some other operating systems.

5 Troubleshooting

All parts of LILO display some messages that can be used to diagnose problems.

5.1 Map installer warnings and errors

Most messages of the map installer (/sbin/lilo) should be self-explanatory. Some messages that indicate common errors are listed below. They are grouped into fatal errors and warnings (non-fatal errors).

5.1.1 Fatal errors

Boot sector of device name doesn't have a boot signature
Boot sector of device name doesn't have a LILO signature

The sector from which LILO should be uninstalled doesn’t appear to be a LILO boot sector.

Can't put the boot sector on logical partition number

An attempt has been made to put LILO’s boot sector on the current root file system partition which is on a logical partition. This usually doesn’t have the desired effect, because common MBRs can only boot primary partitions. This check can be bypassed by explicitly specifying the boot partition with the -b option or by setting the configuration variable boot.

Checksum error

The descriptor table of the map file has an invalid checksum. Refresh the map file immediately !

Device 0xnumber:  Configured as inaccessible.

There is a disk section entry indicating that the device is inaccessible from the BIOS. You should check carefully that all files LILO tries to access when booting are on the right device.

Device 0xnumber:  Got bad geometry sec/hd/cyl

The device driver for your SCSI controller does not support geometry detection. You have to specify the geometry explicitly (see section 3.4).

Device 0xnumber:  Invalid partition table, entry number

The 3D and linear addresses of the first sector of the specified partition don’t correspond. This is typically caused by partitioning a disk with a program that doesn’t align partitions to tracks and later using PC/MS-DOS or OS/2 on that disk. LILO can attempt to correct the problem, see page 22.

Device 0xnumber:  Partition type 0xnumber does not seem suitable for a LILO boot sector

The location where the LILO boot sector should be placed does not seem to be suitable for that. (See also also section 1.1). You should either adjust the partition type to reflect the actual use or put the boot sector on a different partition. This consistency check only yields a warning (i.e. LILO continues) if the option ignore-table is set.

device name is not a valid partition device

The specified device is either not a device at all, a whole disk, or a partition on a different disk than the one in whose section its entry appears.

device name is not a whole disk device

Only the geometry of whole disks (e.g. /dev/hda, /dev/sdb, etc.) can be redefined when using disk sections.

DISKTAB and DISK are mutually exclusive

You cannot use a disktab file and disk geometry definitions in the configuration file at the same time. Maybe /etc/disktab was accidentally used, because that’s the default for backward-compatibility. You should delete /etc/disktab after completing the transition to disk sections.

Duplicate entry in partition table

A partition table entry appears twice. The partition table has to be fixed with fdisk.

Duplicate geometry definition for device name

A disk or partition geometry definition entry for the same device appears twice in the configuration file. Note that you mustn’t write a partition section for the whole disk — its start sector is always the first sector of the disk.

First sector of device doesn't have a valid boot signature

The first sector of the specified device does not appear to be a valid boot sector. You might have confused the device name.^{30}

geo comp addr:  Cylinder number beyond end of media (number)

A file block appears to be located beyond the last cylinder of the disk. This probably indicates an error in the disk geometry specification (see section 3.4) or a file system corruption.

geo comp addr:  Cylinder number is too big (number > 1023)

Blocks of a file are located beyond the 1024th cylinder of a hard disk. LILO can’t access such files, because the BIOS limits cylinder numbers to the range 0…1023. Try moving the file to a different place, preferably a partition that is entirely within the first 1024 cylinders of the disk.

Hole found in map file (location)

The map installer is confused about the disk organization. Please report this error.

item doesn't have a valid LILO signature

The specified item has been located, but is not part of LILO.

item has an invalid stage code (number)

The specified item has probably been corrupted. Try re-building LILO.

item is version number.  Expecting versionnumber.

The specified entity is either too old or too new. Make sure all parts of LILO (map installer, boot loaders and chain loaders) are from the same distribution. ^{31}

Kernel name is too big

The kernel image (without the setup code) is bigger than 512 kbytes. LILO would overwrite itself when trying to load such a kernel. Try removing some unused drivers and compiling the kernel again. With recent (compressed) kernels, the reason for this error is most likely that the kernel image is damaged or that it contains trailing “junk”, e.g. as the result of copying an entire boot floppy to the hard disk.

LOCK and FALLBACK are mutually exclusive

Since lock and fallback both change the default command line, they can’t be reasonably used together.

Map path is not a regular file.

This is probably the result of an attempt to omit writing a map file, e.g. with -m /dev/null. The -t option should be used to accomplish this.

Must specify SECTORS and HEADS together

It is assumed that disks with a “strange” number of sectors will also have a “strange” number of heads. Therefore, it’s all or nothing.

No geometry variables allowed if INACCESSIBLE

If a device is configured as inaccessible (see section 3.4.2), its disk section must not contain any geometry variables.

30Because different partition programs may display the partitions in a different order, it is possible that what you think is your first partition isn’t /dev/hda1, etc. A good method to verify the content of a partition is to try to mount it.

31The expected version number may be different from the version number of the LILO package, because file version numbers are only increased when the file formats change.

No image image is defined

The command line specified either with the -R option or with fallback does not contain the name of a valid image. Note that optional images which have not been included in the map file are not considered as valid.

Partition entry not found

The partition from which an other operating system should be booted isn’t listed in the specified partition table. This either means that an incorrect partition table has been specified or that you’re trying to boot from a logical partition. The latter usually doesn’t work. You can bypass this check by omitting the partition table specification (e.g. omitting the variable table).

Single-key clash:  "name" vs.  "name"

The specified image labels or aliases conflict because one of them is a single character and has the single-key option set, and the other name begins with that character.

Sorry, don't know how to handle device number

LILO uses files that are located on a device for which there is no easy way to determine the disk geometry. Such devices have to be explicitly described, see section 3.4.

This LILO is compiled READONLY and doesn't support . . .

If LILO is not allowed to write to the disk at boot time (see section 4.1.5), options like lock and fallback are unavailable.

This LILO is compiled without REWRITE TABLE and doesn't support . . .

If LILO is not allowed to rewrite partition tables at boot time (see section 3.5), options like activate and set (in a change section) are unavailable. You may also get this error if LILO is compiled with

READONLY enabled.
Timestamp in boot sector of device differs from date of file

The backup copy of the boot sector does not appear to be an ancestor of the current boot sector. If you are absolutely sure that the boot sector is indeed correct, you can bypass this check by using -U instead of -u.

Trying to map files from unnamed device 0xnumber (NFS ?)

This is probably the same problem as described below, only with the root file system residing on NFS.

Trying to map files from your RAM disk.  Please check -r option or ROOT environment variable.

Most likely, you or some installation script is trying to invoke LILO in a way that some of the files is has to access reside on the RAM disk. Normally, the ROOT environment variable should be set to the mount point of the effective root device if installing LILO with a different root directory. See also sections 3.1.2 and 4.1.4.

VGA mode presetting is not supported by your kernel.

Your kernel sources appear to be very old (’93 ?). LILO may work on your system if you remove the vga option.

write item: error reason The disk is probably full or mounted read-only.

5.1.2 Warnings

Messages labeled with “Warning” can be turned off with the nowarn option.

FIGETBSZ file name: error reason The map installer is unable to determine the block size of a file system. It assumes a block size of two sectors (1kB).

Ignoring entry 'variable name'

The command-line option corresponding to the specified variable is set. Therefore, the configuration file entry is ignored.

Setting DELAY to 20 (2 seconds)

Because accidentally booting the wrong kernel or operating system may be very inconvenient on systems that are not run from a local display, the minimum delay is two seconds if the serial variable is set.

(temp) item: error reason Deleting a temporary file has failed for the specified reason.

Warning:  BIOS drive 0xnumber may not be accessible

Because most BIOS versions only support two floppies and two hard disks, files located on additional disks may be inaccessible. This warning indicates that some kernels or even the whole system may be unbootable.

Warning:  COMPACT may conflict with LINEAR on some systems

Please see section 5.3 for a description of this problem.

Warning: config file should be owned by root

In order to prevent users from compromising system integrity, the configuration file should be owned by root and write access for all other users should be disabled.

Warning: config file should be readable only for root if using PASSWORD

Users should not be allowed to read the configuration file when using the password option, because then, it contains unencrypted passwords.

Warning: config file should be writable only for root
     See “Warning: config file should be owned by root”.
Warning:  device 0xnumber exceeds 1024 cylinder limit

A disk or partition exceeds the 1024 cylinder limit imposed by the BIOS. This may result in a fatal error in the current installation run or in later installation runs. See “geo comp addr: Cylinder

     number is too big (number > 1023)” for details.
Warning: device is not on the first disk

The specified partition is probably not on the first disk. LILO’s boot sector can only be booted from the first disk unless some special boot manager is used.

WARNING: The system is unbootable !

One of the last installation steps has failed. This warning is typically followed by a fatal error describing the problem.

5.2 Boot loader messages

The boot loader generates three types of messages: progress and error messages while it is loading, messages indicating disk access errors, and error messages in response to invalid command-line input. Since messages of the latter type are usually self-explanatory, only the two other categories are explained.

5.2.1 LILO start message

When LILO loads itself, it displays the word “LILO”. Each letter is printed before or after performing some specific action. If LILO fails at some point, the letters printed so far can be used to identify the problem. This is described in more detail in the technical overview.

Note that some hex digits may be inserted after the first “L” if a transient disk problem occurs. Unless LILO stops at that point, generating an endless stream of error codes, such hex digits do not indicate a severe problem.

(nothing) No part of LILO has been loaded. LILO either isn’t installed or the partition on which its boot sector is located isn’t active.

L error … The first stage boot loader has been loaded and started, but it can’t load the second stage boot loader. The two-digit error codes indicate the type of problem. (See also section 5.2.2.) This condition usually indicates a media failure or a geometry mismatch (e.g. bad disk parameters, see section 3.4).

LI The first stage boot loader was able to load the second stage boot loader, but has failed to execute it. This can either be caused by a geometry mismatch or by moving /boot/boot.b without running the map installer.

LIL The second stage boot loader has been started, but it can’t load the descriptor table from the map file. This is typically caused by a media failure or by a geometry mismatch.

LIL? The second stage boot loader has been loaded at an incorrect address. This is typically caused by a subtle geometry mismatch or by moving /boot/boot.b without running the map installer.

LIL- The descriptor table is corrupt. This can either be caused by a geometry mismatch or by moving /boot/map without running the map installer.

LILO All parts of LILO have been successfully loaded.

5.2.2 Disk error codes

If the BIOS signals an error when LILO is trying to load a boot image, the respective error code is displayed. The following BIOS error codes are known:

0x00 “Internal error”. This code is generated by the sector read routine of the LILO boot loader whenever an internal inconsistency is detected. This might be caused by corrupt files. Try re-building the map file. Another possible cause for this error are attempts to access cylinders beyond 1024 while using the linear option. See section 1.3.1 for more details and for how to solve the problem.

0x01 “Illegal command”. This shouldn’t happen, but if it does, it may indicate an attempt to access a disk which is not supported by the BIOS. See also “Warning: BIOS drive 0xnumber may not be accessible” in section 5.1.2.

0x02 “Address mark not found”. This usually indicates a media problem. Try again several times.

0x03 “Write-protected disk”. This should only occur on write operations.

0x04 “Sector not found”. This typically indicates a geometry mismatch. If you’re booting a raw-written disk image, verify whether it was created for disks with the same geometry as the one you’re using. If you’re booting from a SCSI disk or a large IDE disk, you should check, whether LILO has obtained correct geometry data from the kernel or whether the geometry definition corresponds to the real disk geometry. (See section 3.4.) Removing compact may help too. So may adding linear.

0x06 “Change line active”. This should be a transient error. Try booting a second time.

0x07 “Invalid initialization”. The BIOS failed to properly initialize the disk controller. You should control the BIOS setup parameters. A warm boot might help too.

0x08 “DMA overrun”. This shouldn’t happen. Try booting again.

0x09 “DMA attempt across 64k boundary”. This shouldn’t happen. Try omitting the compact option.

0x0C “Invalid media”. This shouldn’t happen and might be caused by a media error. Try booting again.

0x10 “CRC error”. A media error has been detected. Try booting several times, running the map installer a second time (to put the map file at some other physical location or to write “good data” over the bad spot), mapping out the bad sectors/tracks and, if all else fails, replacing the media.

0x11 “ECC correction successful”. A read error occurred, but was corrected. LILO does not recognize this condition and aborts the load process anyway. A second load attempt should succeed.

0x20 “Controller error”. This shouldn’t happen.

0x40 “Seek failure”. This might be a media problem. Try booting again.

0x80 “Disk timeout”. The disk or the drive isn’t ready. Either the media is bad or the disk isn’t spinning. If you’re booting from a floppy, you might not have closed the drive door. Otherwise, trying to boot again might help.

0xBB “BIOS error”. This shouldn’t happen. Try booting again. If the problem persists, removing the compact option or adding/removing linear might help.

If the error occurred during a write operation, the error code (two hex digits) is prefixed with a “W”. Although write errors don’t affect the boot process, they might indicate a severe problem, because they usually imply that LILO has tried to write to an invalid location. If spurious write errors occur on a system, it might be a good idea to configure LILO to run read-only (see section 4.1.5).

Generally, invalid geometry and attempts to use more than two disks without a very modern BIOS may yield misleading error codes. Please check carefully if /sbin/lilo doesn’t emit any warnings. Then try using the linear option (see section 3.2.2).

5.3 Other problems

This section contains a collection of less common problems that have been observed. See also section 4.4 for general remarks on using LILO with other operating systems. Some of the problems are obscure and so are the work-arounds.

• If LILO doesn’t go away even if you erase its files, format your Linux partition, etc., you’ve probably installed LILO as your MBR and you’ve forgotten to deinstall it before deleting its files. See section 4.3 for what you can do now.

• For yet unknown reasons, LILO may fail on some systems with AMI BIOS if the “Hard Disk Type 47 RAM area” is set to “0:300” instead of “DOS 1K”.

• Some disk controller BIOSes perform disk geometry/address translations that are incompatible with the way the device’s geometry is seen from Linux, i.e. without going through the BIOS. Particularly, large IDE disks and some PCI SCSI controllers appear to have this problem. In such cases, either the translated geometry has to be specified in a disk section or the sector address translation can be deferred by using the linear option. In a setup where floppies are not normally used for booting, the linear approach should be preferred, because this avoids the risk of specifying incorrect numbers.

• OS/2 is said to be bootable from a logical partition with LILO acting as the primary boot selector if LILO is installed on the MBR, the OS/2 BootManager is on an active primary partition and LILO boots BootManager. Putting LILO on an extended partition instead is said to crash the OS/2 FDISK in this scenario.

Note that booting LILO from BootManager (so BootManager is the primary selector) or booting OS/2 directly from a primary partition (without BootManager) should generally work. See also section 4.4.

• Windows NT is reported to be bootable with LILO when LILO acts as the MBR and the Windows NT boot loader is on the DOS partition. However, NT’s disk manager complains about LILO’s MBR when trying to edit the partition table.

• Some PC UNIX systems (SCO and Unixware have been reported to exhibit this problem) depend on their partition being active. See section 3.5 for how this can be accomplished.

• Future Domain TMC-1680 adapters with the BIOS versions 3.4 and 3.5 assign BIOS device numbers in the wrong order, e.g. on a two-disk system, /dev/sda becomes 0x81 and /dev/sdb becomes 0x80. This can be fixed with the following disk section:

disk=/dev/sda bios=0x81 disk=/dev/sdb bios=0x80

Note that this is only valid for a two-disk system. In three-disk systems, /dev/sdc would become 0x80, etc. Also, single-disk systems don’t have this problem (and the “fix” would break them).

• Some BIOSes don’t properly recognize disks with an unusual partition table (e.g. without any partition marked active) and refuse to boot from them. This can also affect the second hard disk and the problem may only occur if the system is booted in a particular way (e.g. only after a cold boot).

• On some systems, using linear and compact together leads to a boot failure. The exact circumstances under which this happens are still unknown.